Synovec Co. is growing quickly. Dividends are expected to grow at a rate of 17 percent for the next 3 years, with the growth rate falling off to a constant 5 percent thereafter. If the required return is 10 percent and the company just paid a $3.70 dividend. what is the current share price?

Multiple Choice $103.95 $101.62 $108.19 $97.57 $106.07
Answer #1

D1=(3.7*1.17)=4.329

D2=(4.329*1.17)=5.06493

D3=(5.06493*1.17)=5.9259681

Value after year 3=(D3*Growth rate)/(Required rate-Growth rate)

=(5.9259681*1.05)/(0.1-0.05)

=124.44533(Approx).

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.329/1.1+5.06493/1.1^2+5.9259681/1.1^3+124.44533/1.1^3

=$106.07(Approx).